OpenStudy (anonymous):

A garden table and a bench cost 787 combined. The garden table costs 37 more than the bench. What is the cost of the bench?

OpenStudy (madhu.mukherjee.946):

let the cost of the bench be x so garden table=37+x so 2x+37=787 2x=750 x=375

OpenStudy (anonymous):

can you elaborate a little more please but thank you for you help

OpenStudy (madhu.mukherjee.946):

where's my medals for all questions of yours which i have answered

OpenStudy (madhu.mukherjee.946):

say which part u did not understand

OpenStudy (anonymous):

im sorry you deserve medals how do i give you them ? @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

let the cost of the bench be x so the cost of garden table=37+x(cost of garden table is 37 more than the bench) so 2x+37=787 2x=787-37 2x=750 x=375

OpenStudy (madhu.mukherjee.946):

now you understand

OpenStudy (madhu.mukherjee.946):

go to your previous questions and tag me